# A Survey on Data Selection for Language Models

### 26 Feb 2024 v1

@@ 1 Introduction
One often used sense of optimality in data selection is with regard to a model’s performance. In this work, we adopt the commonly held view that, at their core, machine learning models are a method for modeling statistical patterns in data and, from the probabilistic viewpoint, the optimal dataset is that which most closely matches the distribution under which the model will be evaluated (Murphy, 2012). While the probabilistic viewpoint presents a common view of how to select data that improves model performance, this is not the only goal of data selection methods. Data selection methods can reduce costs (by reducing dataset size (Ortiz Suárez et al., 2019; Schreiber et al., 2020; Brown et al., 2020; Lee et al., 2022a; Sorscher et al., 2022)), ensure the integrity of evaluation metrics (by removing data that is suspected to be from the evaluation data (Rae et al., 2022; Marone & Durme, 2023; Oren et al., 2023)), and reducing undesirable behaviors (such as bias and toxicity (Dodge et al., 2021; Welbl et al., 2021; Luccioni & Viviano, 2021; Longpre et al., 2023c)).
Data selection has recently become especially important in the context of large language models. Language models can undergo multiple stages of training (pretraining (Peters et al., 2018; Radford & Narasimhan, 2018; Devlin et al., 2019; Raffel et al., 2020; Touvron et al., 2023a), instruction-tuning (Mishra et al., 2021; Sanh et al., 2022; Longpre et al., 2023a; Muennighoff et al., 2024), alignment (Ziegler et al., 2020; Bai et al., 2022b; Ouyang et al., 2022; Rafailov et al., 2023), etc.), and data selection plays an important role in each stage. However, as the objectives of training differ across each stage, the goals of data selection also vary accordingly. For example, language models are commonly pretrained on large corpora of text from a variety of sources. Among these sources is the internet, where estimates suggest that there are around 250 billion webpages. These webpages amount to around 11 petabytes of data, collected from internet scraping efforts since 2008, with an additional 3-5 billion new web pages being crawled monthly. [^1] According to https://commoncrawl.org/ accessed on 11/02/23 (MM/DD/YY). Due to the massive size of pretraining corpora, a common goal of data selection during pretraining is to remove significant quantities of data through a series of filters (Conneau & Lample, 2019; Raffel et al., 2020; Wenzek et al., 2020; Gao et al., 2020; Rae et al., 2022; Lee et al., 2022a) that aim to only retain data that is deemed “high-quality”. In contrast with pretraining, one form of data selection for fine-tuning a model on a target task is to select additional auxiliary samples that will be most beneficial as additional learning signals for the target task (Albalak et al., 2023b; Ivison et al., 2023a).
@@ 2.1 Background and Motivation of Data Selection
The dataset distribution is also a crucial component of data selection as models can struggle with out-of-distribution generalization and the dataset distribution specifies what data is in-distribution vs. out-of-distribution (Shi et al., 2023b). Specifically, a dense region of the data distribution generally suggests that a model will perform well on unseen data from that region (e.g. similar distribution of tokens). With this in mind, increasing the density of data around desirable regions while reducing density around undesirable regions generally leads to a model that performs well in the desired settings.
@@ 3 Data Selection for Pretraining
Figure 4: An overview of the data filtering pipeline for pretraining. Each filtering component is described in Section 3, and depicts common filters used for preprocessing text data. Note that different works employ different filters, at different stages, and do not necessarily adhere to the order conveyed here.
@@ 3.1 Language Filtering
Language Filtering Summary & Ideas
Language filtering is an important first step of data selection where documents are selected to include only the desired languages. Classifier-based methods are the norm when creating both english-only and multilingual datasets. For lower-resourced languages, URL-based methods are also useful for classifying the language used in multilingual datasets and can sometimes can be more reliable than classifiers. Filtering for code languages is very simple, most works simply using the file extension. [Uncaptioned image] One interesting direction for future research is developing methods that find code data that within documents of natural language. To the best of our knowledge, no such methods currently exist, but a code-language identifier could be used in domains with interleaved code and natural language to find meaningful quantities of new code data.
@@ 3.2 Heuristic Approaches
The current known best practices for heuristics have not changed significantly in the past few years. Soldaini et al. (2024) recently use a combination of the filters from MassiveText (Rae et al., 2022) and c4 (Raffel et al., 2020), while Penedo et al. (2023) also follow the rules defined for MassiveText (Rae et al., 2022). However, it is unclear whether the lack of progress in common heuristics is due to a lack of exploration in the space (due to limited resources and studies), or if the known heuristics truly are optimal. To the best of our knowledge, there have not been studies specifically comparing the many possible heuristics in a systematic study.
Heuristic Approaches Summary & Ideas
Heuristics are a very efficient method for removing large quantities of data, but lack finesse. Most common heuristic utility functions fall into 5 categories: item count, repetition count, existence, ratio, and statistics. Heuristics should be designed specifically for the domain of interest (e.g. heuristics for code will not work for natural language). Heuristics can also be used to improve data point characteristics (e.g. removing HTML from a web document, or specific lines within a document). [Uncaptioned image] Because heuristics lack finesse, it may be worth exploring the use of stochastic selection mechanisms to allow higher quantities of data through the filter (with the hope that a filter farther down the pipeline will remove data that is actually undesirable). [Uncaptioned image] Validating the quality of a heuristic filter is a very slow process, requiring either manual inspection, or training and evaluating a model. Developing methods that can directly measure the data itself could significantly improve iteration time.
@@ 3.3 Data Quality
Training on the highest quality data can lead to stronger performance (Du et al., 2022). However, “high-quality” is not a well defined term with respect to language model pretraining data. Additionally, there is no one-size-fits-all characterization as, depending on the use case, “high-quality” data can vary drastically (Longpre et al., 2023c). Additionally, implementations of toxicity filters and quality filters can further marginalize communities whose text is not commonly considered “high quality” (Xu et al., 2021; Longpre et al., 2023c). In this work, we narrow the use of the phrase “high-quality” to a common use of the phrase referring to: “data that is known to be written by humans, and has likely gone through an editing process” Grave et al. (2018); Gao et al. (2020); Brown et al. (2020); Chowdhery et al. (2022). Some data domains that fall under the “high-quality” category are Wikipedia, books, patents, and peer-reviewed journal articles.
@@
An important consideration when performing quality filtering is that certain reference corpora may be biased towards or away from certain demographics, dialects, and socialects (Rae et al., 2022). Additionally, Gururangan et al. (2022) find that a quality filter trained on Wikipedia, books, and newswire (a replication of the quality filter from Brown et al. (2020)), evaluated on high school newspaper articles, prefers articles from high schools in wealthier, higher-educated, and urban areas. They demonstrate that even though quality filtering is often assumed to be fairly neutral, it carries with it an implied judgment of preferred values. Thus, quality filtering should be done with care to ensure that biases are avoided and the distribution of data sufficiently covers the desired dialects and demographics. Filtering text for quality, while avoiding the introduction of biases is an important direction for future research.
It is still an open question whether using quality filters is a beneficial practice. Some recent works that have entirely foregone using a quality filter include MassiveText (Rae et al., 2022), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024). Specifically, Penedo et al. (2023) elect not to perform any quality filtering and to spend more compute on a stringent deduplication instead. It is not fully understood when quality filtering is beneficial and when it is not required, requiring further research.
Data Quality Summary & Ideas
Data quality methods aim to select data points that are similar to data which is of known “high-quality”. Classifier- and perplexity-based quality filtering are the existing approaches, with classifier-based filtering being the more popular method. For both classifier- and perplexity-based methods, it is important to be careful when training the classifier because this can introduce biases (e.g. against lower-income, less-educated populations). [Uncaptioned image] Exactly which situations quality filtering is beneficial is still undecided, as recent works have trained performant models without using quality filters at all. This is an area that requires further study.
@@ 3.4 Domain-Specific Selection
Domain-Specific Selection Summary & Ideas
Domain-specific selection aims to find data most similar to a specific domain (e.g. medicine or law). Moore-Lewis style methods (requiring one in-domain model and one general purpose model) have stood the test of time and are still the most popular methods for domain-specific filtering. [Uncaptioned image] Datamodels and other methods based on data attribution and valuation can be explored as competitive methods, but generally have higher computational requirements. [Uncaptioned image] Perplexity-based methods may be another direction to explore. They require only an in-domain model, which could improve efficiency. It’s possible that they would not perform as well as Moore-Lewis based methods, but because they use less compute they could be used in multiple rounds of selection, as in bootstrapping.
@@ 3.5 Data Deduplication
Datasets originating from internet dumps often abound with duplicate and near duplicate documents (Elazar et al., 2023; Magnusson et al., 2023). A dataset with duplicated data, or near duplicates, increases the distribution density around those areas. In cases where these high-density points are of high value, keeping near duplicates can be beneficial. However, for pretraining where the exact evaluation distribution is unknown, it is generally preferred to remove duplicates so that the training distribution provides greater coverage with less redundancy. Carlini et al. (2023) demonstrate a direct relation between the number of times a data point is duplicated and the degree to which a model will memorize the data point, a relation which increases with model scale. Additionally, filtering out duplicates can reduce the training time of machine learning models (Sorscher et al., 2022) and has been shown to sometimes improve accuracy on downstream tasks (Lee et al., 2022a; Tirumala et al., 2023).
@@
Similarly to exact matching methods, hash-based methods are very commonly used for approximate matching methods. For example, MinHash (Broder, 1997) is an approximate hashing algorithm that excels at finding templated documents (e.g. licenses with only specific entities differing or placeholder SEO text repeated across websites). Modern deduplication methods use other variants of hashing such as MinHashLSH Brown et al. (2020) and SimHash Charikar (2002) that utilize locality sensitive hashing. [^11] See the blogpost by Mou (2023) for implementation details on MinHashLSH The general method behind hashing methods for deduplication is to first group small segments of each documents (often called shingling), then to encode each segment using a large number of hashing functions into features. Next, aggregate the features (e.g. MinHashLSH keeps only the smallest features). Finally, documents can be compared with some method of similarity computation between features (e.g. MinHashLSH uses a Jaccard similarity) and removing documents below some threshold. When creating MassiveTextRae et al. (2022) use MinHash with 13-grams and 450 hash functions and set the jaccard similarity threshold at 0.8, randomly removing either of the documents that are determined to be duplicates. In contrast, Gao et al. (2020) use only 10 hash functions, but Rae et al. (2022) find that the more aggressive deduplication leads to better model performance.
@@
Finally, model-based approximate matching techniques utilize a learned representation of data points to deduplicate similar data. These methods can be very costly, but allow for more flexibility by considering the semantics of text. For example, SemDeDup (Abbas et al., 2023) uses a pretrained 125M OPT model and calculates a representation for each data point in the C4 dataset by using the last layer embedding of the final token in the document. Then, they cluster data points to create groups of rough similarity, followed by computing the pairwise cosine similarity of data points within each cluster. Document De-duplication and Diversification (D4(Tirumala et al., 2023) go even further, using a pipeline of first deduplicating documents with MinHash, then applying SemDeDup (Abbas et al., 2023) and clustering the remaining data using K-means, and finally applying the SSL prototypes method of Sorscher et al. (2022) which removes the “most prototypical” examples in each cluster. The authors demonstrate that deduplication with multiple methods leads to improved performance as each method is designed to remove only some aspect of similarity. One concern with these methods is that they have only been tested on the C4 dataset, which is known to be quite noisy, and have not been demonstrated to be beneficial on cleaner datasets. Prior work has shown that the embedding performance of pretrained decoder language models such as OPT 125M is suboptimal (Muennighoff, 2022). Thus, a promising future direction for this line of work is using better embedding models that are specifically trained for clustering (Muennighoff et al., 2022a; 2024; Xiao et al., 2023).
@@
Data Deduplication Summary & Ideas
Deduplication is a very important step and is often used multiple times within the data selection pipeline. First with a simple URL-based filter, then with more complicated hashing- and model-based methods. Deduplication generally relies one of four methods: URLs, hashing, string metrics, or model representations. Bloom filters are often used as state-of-the-art exact deduplication (even though they can remove non-duplicated data) as they are very space efficient. For a more thorough deduplication, approximate matching methods can be used, where MinHashLSH is currently the most commonly used method. [Uncaptioned image] In settings where data efficiency is the primary goal, model-based deduplication has demonstrated good results. However, these methods are relatively new and need further study to demonstrate their benefits in other cases (e.g. when performance is the primary goal).
@@ 3.6 Filtering Toxic and Explicit Content
One of the most common methods simply filters documents containing words or phrases that match a lexicon. It is often unclear what an appropriate threshold is for this strategy (e.g., whether a single match, or multiple matches, are required to justify the exclusion of an entire document), but many datasets use such methods (Raffel et al., 2020; Xue et al., 2021; OpenAI et al., 2023). Laurençon et al. (2022) go further and adapt such lexicon based on the detected language, to tailor such toxicity filters across multiple languages.
@@
Another common approach is to score documents using a classifier that aims to detect “toxic” documents. Such classifiers are often trained using a simple classifier (e.g., a linear model over bag-of-words embeddings) so it can be applied at scale on a large dataset. Others use an external, blackbox tools for classifying toxicity, e.g., Jigsaw, or Google’s SafeSearch [^13] https://support.google.com/websearch/answer/510 (Rae et al., 2022; Longpre et al., 2023c).
@@
Toxic and Explicit Content Filtering Summary & Ideas
There is a wide variety of toxic and explicit content that are generally undesirable to have in training or evaluation data. Common methods for filtering out toxic and explicit content include blacklisting certain web domains, blacklisting certain words, using a classifier to detect harmful content, using perplexity to detect harmful content, and removing PII with regular expressions or classifiers. [Uncaptioned image] There have been limited studies on the effects of these filters, warranting further research on the impacts of such filtering to understand how they affect the distribution of data and the downstream models behaviors. [Uncaptioned image] A useful direction for research is how to mitigate any negative impacts of this filtering. Or conversely, how can toxic and explicit behaviors be reduced at other stages of model development (e.g. alignment), or possibly even at evaluation time through decoding methods.

### 08 Mar 2024 v2

@@ 1 Introduction
One often used sense of optimality in data selection is with regard to a model’s performance. In this work, we adopt the commonly held view that, at their core, machine learning models are a method for modeling statistical patterns in data and, from the probabilistic viewpoint, the optimal dataset is that which most closely matches the distribution under which the model will be evaluated (Murphy, 2012). While the probabilistic viewpoint presents a common view of how to select data that improves model performance, this is not the only goal of data selection methods. Data selection methods can reduce costs (by reducing dataset size (Ortiz Suárez et al., 2019; Schreiber et al., 2020; Brown et al., 2020; Lee et al., 2022a; Sorscher et al., 2022)), ensure the integrity of evaluation metrics (by removing data that is suspected to be from the evaluation data (Rae et al., 2021; Marone & Van Durme, 2024; Oren et al., 2024)), and reducing undesirable behaviors (such as bias and toxicity (Dodge et al., 2021; Welbl et al., 2021; Luccioni & Viviano, 2021; Longpre et al., 2023c)).
Data selection has recently become especially important in the context of large language models. Language models can undergo multiple stages of training (pretraining (Peters et al., 2018; Radford & Narasimhan, 2018; Devlin et al., 2019; Raffel et al., 2020; Touvron et al., 2023a), instruction-tuning (Mishra et al., 2021; Sanh et al., 2022; Longpre et al., 2023a; Muennighoff et al., 2024), alignment (Ziegler et al., 2019; Bai et al., 2022b; Ouyang et al., 2022; Rafailov et al., 2023), etc.), and data selection plays an important role in each stage. However, as the objectives of training differ across each stage, the goals of data selection also vary accordingly. For example, language models are commonly pretrained on large corpora of text from a variety of sources. Among these sources is the internet, where estimates suggest that there are around 250 billion webpages. These webpages amount to around 11 petabytes of data, collected from internet scraping efforts since 2008, with an additional 3-5 billion new web pages being crawled monthly. [^1] According to https://commoncrawl.org/ accessed on 11/02/23 (MM/DD/YY). Due to the massive size of pretraining corpora, a common goal of data selection during pretraining is to remove significant quantities of data through a series of filters (Conneau & Lample, 2019; Raffel et al., 2020; Wenzek et al., 2020; Gao et al., 2020; Rae et al., 2021; Lee et al., 2022a) that aim to only retain data that is deemed “high-quality”. In contrast with pretraining, one form of data selection for fine-tuning a model on a target task is to select additional auxiliary samples that will be most beneficial as additional learning signals for the target task (Albalak et al., 2023b; Ivison et al., 2023a).
@@ 2.1 Background and Motivation of Data Selection
The dataset distribution is also a crucial component of data selection as models can struggle with out-of-distribution generalization and the dataset distribution specifies what data is in-distribution vs. out-of-distribution (Shi et al., 2023). Specifically, a dense region of the data distribution generally suggests that a model will perform well on unseen data from that region (e.g. similar distribution of tokens). With this in mind, increasing the density of data around desirable regions while reducing density around undesirable regions generally leads to a model that performs well in the desired settings.
@@ 3 Data Selection for Pretraining
Figure 4: An overview of the data filtering pipeline for pretraining. Each filtering component is described in Section 3, and depicts common filters used for preprocessing text data. Note that different works employ different filters, at different stages, and do not necessarily adhere to the order conveyed here. This figure was adopted and modified from Soldaini et al. (2024)
@@ 3.1 Language Filtering
Language filtering is an important first step of data selection where documents are selected to include only the desired languages. Classifier-based methods are the norm when creating both english-only and multilingual datasets. For lower-resourced languages, URL-based methods are also useful for classifying the language used in multilingual datasets and can sometimes can be more reliable than classifiers. Filtering for code languages is very simple, most works simply using the file extension. [Uncaptioned image] One interesting direction for future research is developing methods that find code data within documents of natural language. To the best of our knowledge, no such methods currently exist, but a code-language identifier could be used in domains with interleaved code and natural language to find meaningful quantities of new code data. Language Filtering Summary & Ideas
@@ 3.2 Heuristic Approaches
The current known best practices for heuristics have not changed significantly in the past few years. Soldaini et al. (2024) recently use a combination of the filters from MassiveText (Rae et al., 2021) and c4 (Raffel et al., 2020), while Penedo et al. (2023) also follow the rules defined for MassiveText (Rae et al., 2021). However, it is unclear whether the lack of progress in common heuristics is due to a lack of exploration in the space (due to limited resources and studies), or if the known heuristics truly are optimal. To the best of our knowledge, there have not been studies specifically comparing the many possible heuristics in a systematic study.
Heuristics are a very efficient method for removing large quantities of data, but lack finesse. Most common heuristic utility functions fall into 5 categories: item count, repetition count, existence, ratio, and statistics. Heuristics should be designed specifically for the domain of interest (e.g. heuristics for code will not work for natural language). Heuristics can also be used to improve data point characteristics (e.g. removing HTML from a web document, or specific lines within a document). [Uncaptioned image] Because heuristics lack finesse, it may be worth exploring the use of stochastic selection mechanisms to allow higher quantities of data through the filter (with the hope that a filter farther down the pipeline will remove data that is actually undesirable). [Uncaptioned image] Validating the quality of a heuristic filter is a very slow process, requiring either manual inspection, or training and evaluating a model. Developing methods that can directly measure the data itself could significantly improve iteration time. Heuristic Approaches Summary & Ideas
@@ 3.3 Data Quality
Training on the highest quality data can lead to stronger performance (Du et al., 2022). However, “high-quality” is not a well defined term with respect to language model pretraining data. Additionally, there is no one-size-fits-all characterization as, depending on the use case, “high-quality” data can vary drastically (Longpre et al., 2023c). Additionally, implementations of toxicity filters and quality filters can further marginalize communities whose text is not commonly considered “high quality” (Xu et al., 2021; Longpre et al., 2023c). In this work, we narrow the use of the phrase “high-quality” to a common use of the phrase referring to: data that is known to be written by humans, and has likely gone through an editing process (Grave et al., 2018; Gao et al., 2020; Brown et al., 2020; Chowdhery et al., 2023). Some data domains that fall under the “high-quality” category are Wikipedia, books, patents, and peer-reviewed journal articles.
@@
Recently, Wettig et al. (2024) have explored a very different direction, and create a utility function from the outputs of a language model system (GPT-4) by asking it to rate multiple documents along various dimensions of perceived quality. While their work shows that the model-based ratings can be a good signal for quality, it can be costly to make enough API calls to such systems. Additionally, biases from the system may be carried over into the quality signal.
@@
An important consideration when performing quality filtering is that certain reference corpora may be biased towards or away from certain demographics, dialects, and socialects (Rae et al., 2021). Additionally, Gururangan et al. (2022) find that a quality filter trained on Wikipedia, books, and newswire (a replication of the quality filter from Brown et al. (2020)), evaluated on high school newspaper articles, prefers articles from high schools in wealthier, higher-educated, and urban areas. They demonstrate that even though quality filtering is often assumed to be fairly neutral, it carries with it an implied judgment of preferred values. Thus, quality filtering should be done with care to ensure that biases are avoided and the distribution of data sufficiently covers the desired dialects and demographics. Filtering text for quality, while avoiding the introduction of biases is an important direction for future research.
It is still an open question whether using quality filters is a beneficial practice. Some recent works that have entirely foregone using a quality filter include MassiveText (Rae et al., 2021), RefinedWeb (Penedo et al., 2023), and Dolma (Soldaini et al., 2024). Specifically, Penedo et al. (2023) elect not to perform any quality filtering and to spend more compute on a stringent deduplication instead. It is not fully understood when quality filtering is beneficial and when it is not required, requiring further research.
Data quality methods aim to select data points that are similar to data which is of known “high-quality”. Classifier- and perplexity-based quality filtering are the existing approaches, with classifier-based filtering being the more popular method. For both classifier- and perplexity-based methods, it is important to be careful when training the classifier because this can introduce biases (e.g. against lower-income, less-educated populations). [Uncaptioned image] Exactly which situations quality filtering is beneficial is still undecided, as recent works have trained performant models without using quality filters at all. This is an area that requires further study. Data Quality Summary & Ideas
@@ 3.4 Domain-Specific Selection
Domain-specific selection aims to find data most similar to a specific domain (e.g. medicine or law). Moore-Lewis style methods (requiring one in-domain model and one general purpose model) have stood the test of time and are still the most popular methods for domain-specific filtering. [Uncaptioned image] Datamodels and other methods based on data attribution and valuation can be explored as competitive methods, but generally have higher computational requirements. [Uncaptioned image] Perplexity-based methods may be another direction to explore. They require only an in-domain model, which could improve efficiency. It’s possible that they would not perform as well as Moore-Lewis based methods, but because they use less compute they could be used in multiple rounds of selection, as in bootstrapping. Domain-Specific Selection Summary & Ideas
@@ 3.5 Data Deduplication
Datasets originating from internet dumps often abound with duplicate and near duplicate documents (Elazar et al., 2023; Magnusson et al., 2023). A dataset with duplicated data, or near duplicates, increases the distribution density around those areas. In cases where these high-density points are of high value, keeping near duplicates can be beneficial. However, for pretraining where the exact evaluation distribution is unknown, it is generally preferred to remove duplicates so that the training distribution provides greater coverage with less redundancy. Carlini et al. (2023) demonstrate a direct relation between the number of times a data point is duplicated and the degree to which a model will memorize the data point, a relation which increases with model scale. Additionally, filtering out duplicates can reduce the training time of machine learning models (Sorscher et al., 2022) and has been shown to sometimes improve accuracy on downstream tasks (Lee et al., 2022a; Tirumala et al., 2024).
@@
Similarly to exact matching methods, hash-based methods are very commonly used for approximate matching methods. For example, MinHash (Broder, 1997) is an approximate hashing algorithm that excels at finding templated documents (e.g. licenses with only specific entities differing or placeholder SEO text repeated across websites). Modern deduplication methods use other variants of hashing such as MinHashLSH (Brown et al., 2020) and SimHash (Charikar, 2002) that utilize locality sensitive hashing. [^11] See the blogpost by Mou (2023) for implementation details on MinHashLSH The general method behind hashing methods for deduplication is to first group small segments of each documents (often called shingling), then to encode each segment using a large number of hashing functions into features. Next, aggregate the features (e.g. MinHashLSH keeps only the smallest features). Finally, documents can be compared with some method of similarity computation between features (e.g. MinHashLSH uses a Jaccard similarity) and removing documents below some threshold. When creating MassiveTextRae et al. (2021) use MinHash with 13-grams and 450 hash functions and set the jaccard similarity threshold at 0.8, randomly removing either of the documents that are determined to be duplicates. In contrast, Gao et al. (2020) use only 10 hash functions, but Rae et al. (2021) find that the more aggressive deduplication leads to better model performance.
@@
Finally, model-based approximate matching techniques utilize a learned representation of data points to deduplicate similar data. These methods can be very costly, but allow for more flexibility by considering the semantics of text. For example, SemDeDup (Abbas et al., 2023) uses a pretrained 125M OPT model and calculates a representation for each data point in the C4 dataset by using the last layer embedding of the final token in the document. Then, they cluster data points to create groups of rough similarity, followed by computing the pairwise cosine similarity of data points within each cluster. Document De-duplication and Diversification (D4(Tirumala et al., 2024) go even further, using a pipeline of first deduplicating documents with MinHash, then applying SemDeDup (Abbas et al., 2023) and clustering the remaining data using K-means, and finally applying the SSL prototypes method of Sorscher et al. (2022) which removes the “most prototypical” examples in each cluster. The authors demonstrate that deduplication with multiple methods leads to improved performance as each method is designed to remove only some aspect of similarity. One concern with these methods is that they have only been tested on CommonCrawl data, which is known to be quite noisy, and have not been demonstrated to be beneficial on cleaner datasets. Prior work has shown that the embedding performance of pretrained decoder language models such as OPT 125M is suboptimal (Muennighoff, 2022). Thus, a promising future direction for this line of work is using better embedding models that are specifically trained for clustering (Muennighoff et al., 2022a; 2024; Xiao et al., 2023).
@@
Deduplication is a very important step and is often used multiple times within the data selection pipeline. First with a simple URL-based filter, then with more complicated hashing- and model-based methods. Deduplication generally relies one of four methods: URLs, hashing, string metrics, or model representations. Bloom filters are often used as state-of-the-art exact deduplication (even though they can remove non-duplicated data) as they are very space efficient. For a more thorough deduplication, approximate matching methods can be used, where MinHashLSH is currently the most commonly used method. [Uncaptioned image] In settings where data efficiency is the primary goal, model-based deduplication has demonstrated good results. However, these methods are relatively new and need further study to demonstrate their benefits in other cases (e.g. when performance is the primary goal). Data Deduplication Summary & Ideas
@@ 3.6 Filtering Toxic and Explicit Content
One of the most common methods simply filters documents containing words or phrases that match a lexicon. It is often unclear what an appropriate threshold is for this strategy (e.g., whether a single match, or multiple matches, are required to justify the exclusion of an entire document), but many datasets use such methods (Raffel et al., 2020; Xue et al., 2021; Achiam et al., 2023). Laurençon et al. (2022) go further and adapt such lexicon based on the detected language, to tailor such toxicity filters across multiple languages.
@@
Another common approach is to score documents using a classifier that aims to detect “toxic” documents. Such classifiers are often trained using a simple classifier (e.g., a linear model over bag-of-words embeddings) so it can be applied at scale on a large dataset. Others use an external, blackbox tools for classifying toxicity, e.g., Jigsaw, or Google’s SafeSearch [^13] https://support.google.com/websearch/answer/510 (Rae et al., 2021; Longpre et al., 2023c).
@@
There is a wide variety of toxic and explicit content that are generally undesirable to have in training or evaluation data. Common methods for filtering out toxic and explicit content include blacklisting certain web domains, blacklisting certain words, using a classifier to detect harmful content, using perplexity to detect harmful content, and removing PII with regular expressions or classifiers. [Uncaptioned image] There have been limited studies on the effects of these filters, warranting further research on the impacts of such filtering to understand how they affect the distribution of data and the downstream models behaviors. [Uncaptioned image] A useful direction for research is how to mitigate any negative impacts of this filtering. Or conversely, how can toxic and explicit behaviors be reduced at other stages of model development (e.g. alignment), or possibly even at evaluation time through decoding methods. Toxic and Explicit Content Filtering Summary & Ideas